Find the exact values of the six trig functions of “theta” in standard position for P(6, 2).
Answer by Alan3354(69443)

Find the exact values of the six trig functions of “theta” in standard position for P(6,2).
------------
x = 6
y = 2
---
r = sqrt(x^2 + y^2) = 2sqrt(10)
-------
sin = y/r
cos = x/r
tan = y/x
---
cot = 1/tan
sec = 1/cos
csc = 1/sin